if the vector 2i+3j+8k is perpendicular to the vector 4i-4j+ak then the value of a is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Ravi
 
If both the vectors are perpendicular. Then their dot product should be 0.
Hence
(2i + 3j + 8k). (4i - 4j +ak) = 8 -12 +8a = 0
8a = 4
a = 0.5
 
 
Regards
Arun (askIITians forum expert)